There are two countries, Anihc (country A) and Bapan (country B), with the same production function...

  1. There are two countries, Anihc (country A) and Bapan (country B), with the same production function fk=5k0.5. However, country A has saving rates of 0.2, depreciation rate of 0.2 and population growth of 0.2; while country B has saving rates of 0.1, depreciation rate of 0.15 and population growth of 0.05. Using the Solow model:
    1. Find the steady state capital-labor ratio for each country.
    2. Find the steady state output per worker, and the steady state consumption per worker for each country.
    3. Which country produces the most per capita? Which country consumes the most per capita?
Answer #1

Q) prod funcn flk) = 5k 5 f(R) = 56°5 SA=02, 8=02n=02 SB = .1, 88 = 15., 13.05 a) Cht8 k at Steady state, sy = (n+8) k 81S (o= (1-5)y + CA = (1-SAYA #CBE (1-1)B Congh. •8x 12.5 10 Ang .9 X.125 = 11.25 C) as y is same for both Nation , 80 both have s